OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (nincompoop):

well first you want to solve the equation of a line by using the y-intercept form y = mx + b obtain or solve for your slope (m) then plug one of the coordinates into the point-slope form y-y1 = m(x-x1) once that is done, you can then rearrange your equation into the standard form Ax+By = c A has to be positive and everything has to be in whole numbers

OpenStudy (anonymous):

Thank you ^.^ but my issue is that I can't figure out how to get everything into whole numbers ;-;

OpenStudy (nincompoop):

what are the values you have?

OpenStudy (nincompoop):

or better yet, what is your y-intercept form?

OpenStudy (anonymous):

The only things I was given is the coordinates (7, -3), (4, -8)

OpenStudy (nincompoop):

solve the equation using the y-intercept form first. follow the steps I provided

OpenStudy (anonymous):

without rearanging I got y=1 2/3x+17 1/3

OpenStudy (nincompoop):

assuming that your y-intercept is correct rearrange it so that it is in the form of Ax+By = c then we'll worry about turning them into whole numbers after

OpenStudy (anonymous):

I think it's 1 2/3x +y=17 1/3

OpenStudy (nincompoop):

so what you wanted to do is move the x value into the left side, correct? for you to accomplish that, you need to subtract that same value on both sides of the equation. it will cancel out the x on the right side and leave you an x value on the left

OpenStudy (anonymous):

So that would be either 5/3x=17 1/3-y, or 5/3x=(-y)+17 1/3?

OpenStudy (nincompoop):

?

OpenStudy (nincompoop):

get a pen and paper, write things down

OpenStudy (anonymous):

I am writing things down. I'm sorry, I'm really bad at this. I got 5/3x=17 1/3-y

OpenStudy (nincompoop):

you want the x and y on the left side you said you had y=1 2/3x+17 1/3 so if you want to move the x value into the left side so it sits besides the y, you want to subtract whatever you have as x on both sides - the left and right of the equation -5/3x + y = -5/3x + 5/3x + 52/3

OpenStudy (nincompoop):

does it make sense?

OpenStudy (anonymous):

OH. Oopsies. So would that be the final answer, or do I need to do more to it? And I think it makes sense

OpenStudy (nincompoop):

if you read what I put earlier, the x value has to be POSITIVE and EVERYTHING has to be WHOLE NUMBERS

OpenStudy (anonymous):

I thought fractions werent whole numbers though .-.

OpenStudy (nincompoop):

right now our x is negative and everything is pretty much in improper fraction form but notice how all of our denominator is 3? and since our x is negative we want to multiply it with another negative to make it positive (negative times negative equals positive)

OpenStudy (nincompoop):

4/4 is a fraction it is a whole number since 4/4 = 1

OpenStudy (anonymous):

I'm sorry, I meant fractions like 5/3, and 52/3

OpenStudy (nincompoop):

what about them? those are improper fractions, I converted them to that form (I hope you know how to do and undo that), so that when we make everything a whole number, you will see a nifty trick

OpenStudy (nincompoop):

so what do you think we should multiply everything by so that they all are whole numbers and so that our x becomes positive?

OpenStudy (anonymous):

Yeah, I know hot to undo it. And..would -3/3 work?

OpenStudy (nincompoop):

if I had 9/2 and I want to convert it to whole number, all I need to do is multiply it by 2 9/2 * 2 = 9 if I had 97/5 and want to make it a whole number, all I need to do is multiply it by 5 97/5 * 5 = 97 if I had -67/15 and would like to make it a positive whole number, all I need to do is multiply it by -15 -97/5 * -15 = 97 see the pattern here?

OpenStudy (anonymous):

So multiply by -3? I got 5x-3y=52. Am I doing this right?

OpenStudy (nincompoop):

there you go, you've converted something into standard form of equation of a line

OpenStudy (anonymous):

Thank you! So much c:

OpenStudy (anonymous):

And your help will be too ^.^

OpenStudy (anonymous):

Good night, and thanks again ^.^